# Wasserstein Proximal Coordinate Gradient Algorithms

### v1

@@ Appendix B Proof of Lemmas in the Main Theorems
$\int_{\mathcal{X}_{j}}\big{\|}\nabla V_{j}^{k}+\nabla h_{j}^{\prime}(\widetilde{\rho}_{j}^{k})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\widetilde{\rho}_{j}^{k})\big{\|}^{2}\,{\rm d}\widetilde{\rho}_{j}^{k}$ $=\int_{\mathcal{X}_{j}}\Big{\|}\nabla V_{j}^{k}-\nabla V_{j}^{k-1}+\frac{1}{\tau}\Big{(}T_{\widetilde{\rho}_{j}^{k}}^{\widetilde{\rho}_{j}^{k-1}}-\textrm{Id}-\eta_{j}^{k}\Big{)}\Big{\|}^{2}\,{\rm d}\widetilde{\rho}_{j}^{k}$
@@
By Lemma B.1, $T_{\rho_{j}^{k}}^{\rho_{j}^{k-1}}-\textrm{Id}=\tau\nabla V_{j}^{k-1}+\tau\nabla h_{j}^{\prime}(\rho_{j}^{k})+\tau\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k})$. Therefore, we have
@@
Here, (i) is by Cauchy–Schwarz inequality, and (ii) is by Lemma C.1. Summing the above inequality from $j=1$ to $m$ yields the desired result.
@@
be a function on $\mathcal{X}_{j}$. By Lemma B.1, we have $T_{\rho_{j}^{k}}^{\rho_{j}^{k-1}}-\textrm{Id}=\tau\big{[}\nabla\widetilde{V}_{j}^{k}+\nabla h_{j}^{\prime}(\rho_{j}^{k})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k})\big{]}$. Therefore, we have
@@
Here, (i) is by Cauchy–Schwarz inequality; (ii) is by Lemma C.1.
B.10 Proof of Lemma B.3
Proof.
For any $j\in[m]$ and $k\in\mathbb{N}$, let $I_{j,k}\in[M]$ be the time of the latest update of the $j$-th coordinate in the $k$-th iteration. By definition, we know $\rho_{j}^{k-1,M}=\rho_{j}^{k}=\rho_{j}^{k-1,I_{j,k}}$ for all $j\in[m]$. Since
@@
Here, (i) is by Lemma C.1. In the $k$-th iteration, recall that $j_{l}\in[m]$ be the coordinate updated in the $l$-th update. The goal is to bound (27).
@@
Combining all pieces above yields the desired result. ∎
@@ Appendix C Other technical lemmas
for simplicity. Since $\rho^{\ast}$ is the stationary point of the optimization problem (1) on the Wasserstein space, we know $\nabla V_{j}^{\ast}+h_{j}^{\prime}(\rho_{j}^{\ast})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{\ast})=0$ (e.g., using Lemma B.2). This implies
@@
C.2 Proof of Lemma B.1 (first-order optimality condition)
Proof.
Since $\rho\in\mathscr{P}_{2}^{r}(\mathcal{X})$, by Brenier’s theorem (Theorem 2.12 in [48]), there is a unique optimal map $\nabla\phi$ from $\rho_{\tau}$ to $\rho$ and the Kantorovich potential $\phi$ is convex. Then, by Lemma 2.1 in [17] we know $\phi$ is unique up to a constant. By Proposition 7.17 in [42], we know the first variation of the Wasserstein term is $\frac{\phi}{\tau}$. The first variation of the potential energy term and the self-interaction term are $U$ and $\int W(\cdot,y){\rm d}\rho_{\tau}(y)+\int W(y,\cdot){\rm d}\rho_{\tau}(y)$ respectively (see the first bullet point of Remark 7.13 in [42]). Next, we will show the first variation of the internal energy functional is $h^{\prime}(\rho_{\tau})$.
Let $\tilde{\rho}=c=|\mathcal{X}|^{-1}$ be the constant positive density on $\mathcal{X}$, and let $\rho_{\varepsilon}=(1-\varepsilon)\rho_{\tau}+\varepsilon\tilde{\rho}$. Similar to Lemma 8.6 in [42], we have $C_{U}\varepsilon\geq\varepsilon(\rho_{\tau}-c)h^{\prime}(\rho_{\varepsilon})$ where $C_{U}$ is a positive value depending on $U$. Let
@@
is a constant [$\rho_{\tau}$]-a.e.. Taking the gradient yields the result. ∎
C.3 Proof of Lemma C.1
Lemma C.1.
@@
Proof.
Just notice that
@@
Here, (i) is by Cauchy–Schwarz inequality; (ii) is by the $L$-smoothness of $V$. ∎
@@ C.4 Tensorization of Wasserstein distance
Proof.
By definition, we have
@@
The inequality holds while taking $\pi=(\textrm{Id},T_{\rho}^{\rho^{\prime}})_{\#}\rho$. This implies the desired result. ∎
C.5 Proof of Proposition 3.1
@@ C.6 Lemmas for controlling numerical error
Proof.
It is easy to see that
@@
Similar to the proof in [Proof of Theorem 4, 54], there is a constant $C=C(\alpha,A)>0$ such that
@@
Proof.
Without loss of generality, we assume $\xi_{k}\leq\frac{1}{k^{\alpha}}$. Otherwise, we have
@@
Since $\alpha<1$, we just need $AK^{2}-(B+2^{\alpha}\alpha)K-C\geq 0$. By induction, we know such $K$ exists. ∎
@@ C.7 Numerical methods in FA approach
For $\mathcal{H}(\rho^{k})=\int\rho^{k}\log\rho^{k}$, we can apply the argument in [35] by noting that
@@
Since the first term does not depend on $T$, the above discussions yield the approximation in Example 3.2
@@
where $\widehat{\rho}_{\rm kde}^{k}$ is the kernel density estimation of $\rho^{k}$ by $\{X_{b}^{k}:b\in[B]\}$. This leads to the approximation in Example 3.3.
@@ Appendix D Results for Multi-species Systems
D.1 Proof of Proposition 5.3
Since $\rho^{\ast}=(\rho_{1}^{\ast},\dots,\rho_{m}^{\ast})$ is the minimum of (1), by Lemma B.1, we have
@@ Appendix E Connections Between Different Lipschitz Conditions
Here, (i) is due to Section 3.2 in [50]. Thus, we have $L/L_{c}\leq\sqrt{m-1}$. To see the bounds are tight, consider the examples $V(x_{1},x_{2})=x_{1}^{2}+x_{2}^{2}$ and $V(x_{1},\cdots,x_{n})=(x_{m}+\cdots,x_{m})^{2}$. In the first example, we have $L=0$ and $L_{c}=2$, indicating that $L/L_{c}=0$. In the second example, we have $L=2\sqrt{m-1}$ and $L_{c}=2$, indicating that $L/L_{c}=\sqrt{m-1}$.

### v2

@@ B.8 Proof of Lemma 8
$\quad\,\int_{\mathcal{X}_{j}}\big{\|}\nabla V_{j}^{k}+\nabla h_{j}^{\prime}(\widetilde{\rho}_{j}^{k})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\widetilde{\rho}_{j}^{k})\big{\|}^{2}\,{\rm d}\widetilde{\rho}_{j}^{k}$
$=\int_{\mathcal{X}_{j}}\Big{\|}\nabla V_{j}^{k}-\nabla V_{j}^{k-1}+\frac{1}{\tau}\Big{(}T_{\widetilde{\rho}_{j}^{k}}^{\widetilde{\rho}_{j}^{k-1}}-\textrm{Id}-\eta_{j}^{k}\Big{)}\Big{\|}^{2}\,{\rm d}\widetilde{\rho}_{j}^{k}$
@@ B.9 Proof of Lemma 10
By Lemma 9, $T_{\rho_{j}^{k}}^{\rho_{j}^{k-1}}-\textrm{Id}=\tau\nabla V_{j}^{k-1}+\tau\nabla h_{j}^{\prime}(\rho_{j}^{k})+\tau\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k})$. Therefore, we have
@@
Here, (i) is by Cauchy–Schwarz inequality, and (ii) is by Lemma 12. Summing the above inequality from $j=1$ to $m$ yields the desired result.
@@
be a function on $\mathcal{X}_{j}$. By Lemma 9, we have $T_{\rho_{j}^{k}}^{\rho_{j}^{k-1}}-\textrm{Id}=\tau\big{[}\nabla\widetilde{V}_{j}^{k}+\nabla h_{j}^{\prime}(\rho_{j}^{k})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k})\big{]}$. Therefore, we have
@@
Here, (i) is by Cauchy–Schwarz inequality; (ii) is by Lemma 12.
B.10 Proof of Lemma 11
Proof  For any $j\in[m]$ and $k\in\mathbb{N}$, let $I_{j,k}\in[M]$ be the time of the latest update of the $j$-th coordinate in the $k$-th iteration. By definition, we know $\rho_{j}^{k-1,M}=\rho_{j}^{k}=\rho_{j}^{k-1,I_{j,k}}$ for all $j\in[m]$. Since
@@ B.10 Proof of Lemma 11
Here, (i) is by Lemma 12. In the $k$-th iteration, recall that $j_{l}\in[m]$ be the coordinate updated in the $l$-th update. The goal is to bound (27).
@@
Combining all pieces above yields the desired result.
@@ C.1 Proof of Proposition 8
for simplicity. Since $\rho^{\ast}$ is the stationary point of the optimization problem (1) on the Wasserstein space, we know $\nabla V_{j}^{\ast}+h_{j}^{\prime}(\rho_{j}^{\ast})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{\ast})=0$ (e.g., using Lemma 10). This implies
@@
C.2 Proof of Lemma 9 (first-order optimality condition)
Proof  Since $\rho\in\mathscr{P}_{2}^{r}(\mathcal{X})$, by Brenier’s theorem (Theorem 2.12 in (Villani, 2021)), there is a unique optimal map $\nabla\phi$ from $\rho_{\tau}$ to $\rho$ and the Kantorovich potential $\phi$ is convex. Then, by Lemma 2.1 in (Del Barrio and Loubes, 2019) we know $\phi$ is unique up to a constant. By Proposition 7.17 in (Santambrogio, 2015), we know the first variation of the Wasserstein term is $\frac{\phi}{\tau}$. The first variation of the potential energy term and the self-interaction term are $U$ and $\int W(\cdot,y){\rm d}\rho_{\tau}(y)+\int W(y,\cdot){\rm d}\rho_{\tau}(y)$ respectively (see the first bullet point of Remark 7.13 in (Santambrogio, 2015)). Next, we will show the first variation of the internal energy functional is $h^{\prime}(\rho_{\tau})$.
Let $\tilde{\rho}=c=|\mathcal{X}|^{-1}$ be the constant positive density on $\mathcal{X}$, and let $\rho_{\varepsilon}=(1-\varepsilon)\rho_{\tau}+\varepsilon\tilde{\rho}$. Similar to Lemma 8.6 in (Santambrogio, 2015), we have $C_{U}\varepsilon\geq\varepsilon(\rho_{\tau}-c)h^{\prime}(\rho_{\varepsilon})$ where $C_{U}$ is a positive value depending on $U$. Let
@@ C.2 Proof of Lemma 9 (first-order optimality condition)
is a constant [$\rho_{\tau}$]-a.e.. Taking the gradient yields the result.
C.3 Proof of Lemma 12
Lemma 12.
@@ C.3 Proof of Lemma 12
Proof  Just notice that
@@
Here, (i) is by Cauchy–Schwarz inequality; (ii) is by the $L$-smoothness of $V$.
@@ C.4 Tensorization of Wasserstein Distance
Proof  By definition, we have
@@
The inequality holds while taking $\pi=(\textrm{Id},T_{\rho}^{\rho^{\prime}})_{\#}\rho$. This implies the desired result.
C.5 Proof of Proposition 4
@@ C.6 Lemmas for Controlling Numerical Error
Proof  It is easy to see that
@@
Similar to the proof in (Proof of Theorem 4, Yao et al., 2023), there is a constant $C=C(\alpha,A)>0$ such that
@@
Proof  Without loss of generality, we assume $\xi_{k}\leq\frac{1}{k^{\alpha}}$. Otherwise, we have
@@
Since $\alpha<1$, we just need $AK^{2}-(B+2^{\alpha}\alpha)K-C\geq 0$. By induction, we know such $K$ exists.
@@ C.7 Numerical Methods in FA Approach
For $\mathcal{H}(\rho^{k})=\int\rho^{k}\log\rho^{k}$, we can apply the argument in (Mokrov et al., 2021) by noting that
@@
Since the first term does not depend on $T$, the above discussions yield the approximation in Example 5
@@
where $\widehat{\rho}_{\rm kde}^{k}$ is the kernel density estimation of $\rho^{k}$ by $\{X_{b}^{k}:b\in[B]\}$. This leads to the approximation in Example 6.
@@ Appendix D Results for Multi-species Systems
D.1 Proof of Proposition 25
Since $\rho^{\ast}=(\rho_{1}^{\ast},\dots,\rho_{m}^{\ast})$ is the minimum of (1), by Lemma 9, we have
@@ Appendix E Connections between Different Lipschitz Conditions
Here, (i) is due to Section 3.2 in (Wright, 2015). Thus, we have $L/L_{c}\leq\sqrt{m-1}$. To see the bounds are tight, consider the examples $V(x_{1},x_{2})=x_{1}^{2}+x_{2}^{2}$ and $V(x_{1},\cdots,x_{n})=(x_{m}+\cdots,x_{m})^{2}$. In the first example, we have $L=0$ and $L_{c}=2$, indicating that $L/L_{c}=0$. In the second example, we have $L=2\sqrt{m-1}$ and $L_{c}=2$, indicating that $L/L_{c}=\sqrt{m-1}$.